OpenStudy (marmar10):

Help! prove 1+sec^2 x/tanx = 1+cos^2 x/sinxcosx

OpenStudy (anonymous):

Is this your equation?\[1+\frac{sec^2 x}{tanx} = 1+\frac{cos^2x}{sinxcosx}\]

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

Or perhaps this?\[ \frac{1+sec^2 x}{tanx} = \frac{1+cos^2x}{sinxcosx} \]

OpenStudy (marmar10):

the second one

OpenStudy (anonymous):

Okay let's work on the left side: \[ LHS = \frac{1+sec^2(x)}{tan{x}} \] \[ \frac{1+\frac{1}{cos^2(x)}}{\frac{sinx}{cosx}} \] \[ (1+\frac{1}{cos^2(x)})*{\frac{cosx}{sinx}} \] \[ \frac{cosx+\frac{1}{cosx}}{sinx} \] \[ \frac{cosx+\frac{1}{cosx}}{sinx} * \frac{cosx}{cosx} = \frac{1+cos^2x}{sinxcosx} = RHS \]

OpenStudy (marmar10):

this makes so much sense now thank you so much!!